OpenStudy (anonymous):

How would you simplify this rational expression 3x-12 over x-4

OpenStudy (angela210793):

3x-12=3(x-4) so all that would be=3

OpenStudy (anonymous):

3x-12 = 3(x-4) So cancelling out x-4 from numerator and denominator u get 3 as answer

OpenStudy (anonymous):

Firstly I would factorise 3x - 12 which is 3(x-4) Then I would cancel out the (x-4) in the numerator and denominator leaving an answer.. 3

OpenStudy (anonymous):

Thanks Ange

OpenStudy (anonymous):

thanks btw i love that you are giving me the answer but i would like it explained like the one gianfranco did thanks by the way you are all amazing

OpenStudy (angela210793):

right...hmmm....sorry :):):) Will do that next time :)

OpenStudy (anonymous):

ok -6sx over 4x

OpenStudy (angela210793):

-6x/4x? if this is it then we can simplify x-s ---> -6x/4x=-6/4 then we can simplify with 2---> -3/2

OpenStudy (anonymous):

here is the work sheet

OpenStudy (angela210793):

u have 2 do all these?/???????? :O:O:O which r the ones u can't solve?

OpenStudy (anonymous):

6-13

OpenStudy (angela210793):

ok

OpenStudy (angela210793):

7) -6sx/4x? if this is it then we can simplify x-s ---> -6sx/4x=-6/4 then we can simplify with 2---> -3s/2

OpenStudy (anonymous):

so we dont do anything with the s leave it alone

OpenStudy (angela210793):

8) 10w^4/15w we can simplify a w---->10W^3/15 and yes we can't do anything with s...:)

OpenStudy (anonymous):

so you cancel out a w

OpenStudy (angela210793):

I am soo sooo soo sorry...idk wht's wrong with my connection :(:(:( I'm sorry :( 9).....5x-20/(5x+45) we can factorize and simplify a 5---> 5(x-4)/5(X+9)=(x-4)/(x+9) 10)......7x/(-7x+x^2) we can factorize and simplify a x---> 7x/[x*(-7+x)]=x/(-7+x) 11)2x^2/(x^2-2x^3)=we can factorize and simplify a x^2--->2x^2/x^2*(1-2x)=2/(1-2x) 12)....(x+3)/(x^2-3x-18) we can express x^2-3x-18=(x+3)*(x-6) so we can simplify a (x+3) and we'll have =1/(x-6) 13)....(x^2-1)/(1-x)=(x-1)*(x+1)/(1-x)=(x-1)*(x+1)/-(x-1)=-(x+1) Once again i'm sorry and i really hope u'll understand these :):)